OpenStudy (atlchic):

solve by elimination, 2s-3t=-4 and 7-3t=s. Show your work.

OpenStudy (atlchic):

can u show ur work on the whteboard for me

OpenStudy (anonymous):

2s-3t=-4 s+3t=7 add 3s+0=3 so s=1 plug 1 in for s in either eq and 2(1)-3t=-4 2-3t=-4 -3t=-6 t=2

OpenStudy (atlchic):

brb

hero (hero):

I have to do a better job of reading the instructions.

OpenStudy (atlchic):

k

hero (hero):

I would have gotten the correct answer, but I used substitution rather than elimination.

hero (hero):

My mistake.